# Evolving Agents: Interactive Simulation of Dynamic and Diverse Human Personalities

### v1

@@ 3.2.3. Cognition
Emotion perceives external objective events produced by the Behavior system, generating subjective emotions for each behavior. However, this perception is immediate and fragmented. Real humans are not directly shaped by every single occurrence and emotion; instead, through unique cognitive processes, information is filtered and holistically evaluated and summarized. instead,humans will filter and evaluate information we received.Such process is Interestingly, psychologists also apply computational thinking to understand how humans absorb, comprehend information, and shape their personalities, viewing the human brain as an information processing system that identifies, stores, and ultimately reorganizes and reconstructs its system(varela_embodied_2017, ). Based on this concept, we developed the Memory and Insight processes within the Cognition module.
@@
Memory - Characteristic Information Filter: Memory serves as a universal module in current Agent design, with the Generative Agent architecture introducing the Memory-stream architecture for storing, encoding, and retrieving relevant experiences. Cognitive psychology suggests that people with different personalities tend to notice and acquire information about the objective world differently(oren_dynamic_1991, ), indicating the need for a Memory system tailored to individual traits. Therefore, the Evolving Agents places greater emphasis on differentiated information weight assessment. It includes two types of memory data structures: short-term and long-term memory. Short-term memory stores all executed actions along with their corresponding emotions as temporary data. Long-term memory, on the other hand, archives information that has been filtered and summarized from short-term memory. The Agent selects from its existing short-term memory based on the personality features within its Character Structure. It subjectively describes multiple sets of behaviors and emotions, storing the outcomes in its long-term memory database. In practice, Agents indeed develop long-term memories that align with personal traits, especially personality contradictions. For example, Benjamin, who is contemplating interdisciplinary development, pays attention to information outside his field of study, such as reading a literature book in the library. Isabella, grappling with personal versus societal conflicts, takes great interest in the social philosophy discussions provided by Sophia.
@@
Cognitive psychology’s memory processing theory(minsky_society_1988, ) suggests that even long-term memories fade over time, understood as the gradual blurring of information. The design of long-term memory also incorporates this fading mechanism by setting a capacity limit for the memory pool. Upon reaching this threshold, the Agent abstracts and condenses the earliest memories to blur them which is shown in Figure 10.
@@
Figure 10. The long-term memory mechanism of characters involves extracting information of interest from short-term memory and the blurring of older information.
@@
Insight - Overall Process of the Contextual Incident: Cognitive psychology emphasizes the holism of cognition, where people combine information and relevant contexts to complete an integrated cognitive process(minsky_society_1988, ). However, in our system, the information filtered through the memory module remains a vast amount of fragmented data, making it difficult for agents to integrate them for holistic thinking. Therefore, we designed the Insight submodule within the Cognition module to help simulate this holistic thinking process. What’s more, we aim for Insight to reflect the diversity of individual cognitive processes. Thus, we enable Insight to reference the complete character structure, summarizing and reflecting on all events of the day and the long-term memory. In practical tests, we can intuitively perceive that each agent has distinctly unique insights into daily events.
@@
For example, on the first day, Benjamin realizes that becoming a well-rounded individual requires not only technical breakthroughs but also efforts in humanitarian care and social skills. On the second day, Benjamin sees the potential for integrating technology with the humanities, beginning to incorporate a humanitarian perspective into his technical explorations, seeking a harmonious balance between the two. The procedure of such is shown in Figure 11.
@@ 3.2.4. Character Growth
Chain-like Update Process: After processing external events through the Emotion and Cognition modules, the foundational materials for Agent evolution are prepared. The Character Growth module is responsible for simulating the personality update of agents, transforming daily evolutions into comprehensive descriptions updated within the agent’s Character Structure. This structure doesn’t involve complex algorithmic logic but demands high precision in prompt engineering to accurately depict personality changes. We designed the Character Growth process in a chain-like pattern after extensive testing and consideration of the interrelations within the Character Structure, dividing it into four sequential sub-processes based on the Character Structure dimensions, excluding Basic Information: State Modify, Feature Modify, Contradiction Modify and Preference Modify. In the design of specific prompts, we maintain Character Structure dimensions, continuously summarizing new personality information based on the current state of the agent and filling it into the corresponding dimensions. In the continuous assessment of subsequent development, this growth mechanism has provided us with traceable and observable descriptions of personality changes.
@@ 4. Simulation Platform
Like the sandbox simulation environment proposed in the former work, we create a Sandbox using the Unity engine as shown in Figure 12. In this sandbox Agents will move and exhibit their actions based on the computational results of Evolving Agents architecture, making it easier for observers to interact.
Figure 12. In the sandbox world, users can observe the Agent’s daily behavioral activities and inner thoughts.
@@ 4.1.1. Three Level Based Environment
Describing information contained in the environment to the large language model is vital. Agents need to be aware of the objects in the environment; whether they are interactable and how to interact. When describing information contained in the environment,we offer a novel thought. We design a Building-Place-Goal three level data structure, shown in Figure 15, to describe the environment. This structure enables us to efficiently convey structured environmental information to large language models, thus allowing Agents to fully consider existing environmental information and locations during their planning phase, improving the efficiency of Agents to understand the environment. The direct manifestation of this is that large language models will aim to align their planning objectives as closely as possible with the existing environmental functions and descriptions.
@@ 4.1.2. NPC Action Execute
The daily plans generated by the Agent planning module are read and translated into ScriptableObject-type task description files in Unity. When the game time reaches the next plan, the engine will drive the Agent to move to the target location using the A-star algorithm, and upon arrival, play the corresponding animation. To enhance the decoupling of work between the agent architecture and the sandbox, we have detailed programmatic descriptions of the operational rules in sandbox, such as the Agent’s moving speed and spatial information descriptions. This way, the agent architecture can directly generate Agent performance data that fits the logic of the sandbox (for example, if the Sandbox screen running time from the dormitory to the college is 15 minutes, then the agent architecture will take this 15-minute moving time into consideration in advance when executing). This approach not only ensures the decoupling level during the team development process, but also enables user to design and iterate environment setting when art resource for sandbox is not available.
@@ 4.2.1. Environment Editing
Inspired by behavioral psychology, Evolving Agents is also envisioned as a research tool for the field. In behavioral psychology, external stimuli is one of the primary variables in personality development or change. In this potential demand scenario, we hope the system will allow users to easily modify the environmental settings of the system to meet their needs in the future. Environmental updates in the game interface require corresponding art resources, but it is impractical to expect users to provide these resources. However, as mentioned in section 4.2, agent architecture and sandbox are highly decoupled. Users can edit the environment and obtain simulation results without waiting for the art scenes to update. To lower the learning threshold for user editing, we have developed a tabular operation experience for environmental editing. Users can directly control environmental variables by adjusting an online csv table, thereby achieving adjustments to the living space of the Agent.
Figure 14. Users can directly update environmental setting by adjusting an online csv table
@@ 4.2.2. Chat With Agent
Although the current research theme does not encourage users to provide artificial stimuli beyond the simulation process for Agents, we also build relevant user interaction experiences in the Simulative System. We aim for interactions from users to be consistent with other Agents from the perspective of an Agent. Therefore, users can choose to initiate a conversation with any Agent that is not currently engaged in a dialogue during the system’s operation. This conversation will also be recorded and summarized in the Agent’s own Dialog-memory, contributing to the Agent’s cognitive process.
@@ 5. OBJECTIVE DATA VERIFICATION
The evaluation of believability and perceivability mentioned above relies on human users’ observation and subjective assessment. However, before recruiting users for experimentation, we can first analyze the operational logs of the Agents to verify changes in personality and behavior patterns that have occurred during the simulation from a machine metrics standpoint (like alpha and beta testing in software development processes).
We implemented two types of objective data validations, using the Big Five personality assessment and behavior difference analysis respectively. By comparing with the ablation architecture, we objectively verified that the Evolving Agents indeed exhibited personality evolution and corresponding behavioral changes during the simulation process.
@@ 5.1. Big Five Personality Traits Assessment
We aim for Agents to define their current personality based on objective scale questionnaires throughout the simulation process and find the objective evidence of personality changes in Evolving Agents by comparing results over time. The character structure of Evolving Agents is represented based on the Big Five personality traits. Therefore, we refer to existing Big Five personality analysis questionnaires and equip Evolving Agents with the ability to fill out questionnaires, enabling them to continuously undertake Big Five Personality Inventory (BFI-44) assessments(john_big_1999, )and to output the analysis results of the Big Five personality traits during the simulation process. In the specific implementation of enabling Evolving Agents to fill out questionnaires, we discover that even when we constrain the Agent to the same state, the results of each BFI-44 assessment vary significantly. The large language models lack a clear standard for reference and comparison. As a Result, we design a process for filling out the same set of questions across multiple days in parallel, asking the model to focus on the differences from day to day for a comprehensive questionnaire completion which effectively resolves the issue of lacking a clear standard and reference for questionnaire completion across sessions.
Contrasted with the complete Evolving Agents architecture, we have an ablated architecture that lacks the Insight, and Character Growth components of the Personality system. The reason for this ablation is that this version of the Evolving Agents architecture is somewhat similar to previous Agent architectures, including the basic processes of Plan-Action-Memory-Reflect, hence it has certain referential value.
We use three Agents: Isabella, Benjamin, Sophia, for a seven-day simulation within a sandbox, calculating their daily scores across the five dimensions of the Big Five personality traits, as is shown in Figure 16 and Figure 17, with the scoring method referenced. Finally, we determine the difference values between each dimension over two days for each person. Let
@@
be the set of dimensions evaluated. For each dimension $d\in D$, given a series of scores $s_{d}$ defined as shown in below
@@
The Big Five personality change states of Isabella, Benjamin and Sophia, who have different initial personality settings, over seven days are shown below
@@
We can see that the Agents in the ablated group show no change in their personality assessment results over the seven days, with an average personality difference of 0 between every two days for all three Agents. In contrast, Agents in the Full Architecture clearly exhibit active personality change states. For Benjamin, the $\Delta_{\text{overall}}$ is 4.37; for Isabella, it is 6.52; and for Sophia, it is 6.77. This data supports our claim to a great extent that the Evolving Agents can lead to personality evolution in Agents during the simulation process.
From the trends in changes across various dimensions, we also discover that Agents with different personality exhibit distinct pattern of personality evolution. For example, Benjamin, an introverted and stable computer science student, is slowly adjusting his personality towards his target. In contrast, Sophia, a more creative and active writer, experiences rapid changes in both neuroticism and extraversion dimensions, which intuitively aligns with our understanding of their individual personalities.
However, while displaying divergent changes, the three individuals also show certain similar development directions. Overall, all three have shown a decrease in neuroticism, which might be due to the positive bias in the content of large language models, making it easier for agents to evolve towards more self-consistent and peaceful personality directions.
@@ 5.2. Behavioral Similarity
In the Agent architecture, we limit the goal selection in the planning phase to a finite number of abstract categories. This allows us to conveniently measure the similarity of daily plan by encoding each day’s plan and analyze them based on the Euclidean distance, thereby providing a measure of behavioral differences between days. Given a dataset representing daily goal counts, the Euclidean distance between the behavior on two different days, day $i$ and day $j$, is defined by the formula:
@@
where $E(i,j)$ represents the Euclidean distance between day $i$ and day $j$, $x_{ik}$ is the count of goal $k$ on day $i$, and $x_{jk}$ is the count of the same goal on day $j$. $N$ is the total number of distinct goals observed over all days. This calculation assumes that for any goal not observed on a given day, its count is taken as 0.
Given a distance matrix $D$ of size $n\times n$, where $D_{ij}$ represents the Euclidean distance between the behavior on days $i$ and $j$, the overall activity level, denoted as $A$, is calculated as the average of all upper triangular elements of $D$ excluding the diagonal. This overall activity level is given by the formula:
@@
We use the same ablation control group as in the Big Five personality trait analysis for the same reason. This is the version of Evolving Agents most similar in architectural description to past studies, hence it holds referential value.
We use three Agents: Isabella, Benjamin, Sophia for a seven-day simulation within a Sandbox, calculating the behavioral differences between days. The analysis results of behavioral similarity are as follows:
@@
For Isabella, the activity level in the full architecture is 2.293, while the overall activity level in the ablated architecture is 1.416; for Benjamin, the overall activity level in the full architecture is 2.152, while in the ablated architecture it is 1.742; for Sophia, the overall activity level in the full architecture is 1.949, while in the ablated architecture it is 1.474. It is evident that compared with the ablated Agent architecture, Agents in the full architecture exhibited more active behavioral changes.
When specifically planning behavior in the Plan module, there are no differences in the input data composition between the full architecture and the ablated architecture of Agent planning. The only difference is that the full architecture agents receive evolved personalities every day, whereas the ablated architecture repeats the initial personality settings. This also indirectly verifies that personality differences are one of the reasons affecting changes in behavior patterns.

### v3

@@ 3.2.3. Cognition
Emotion perceives external objective events produced by the Behavior system, generating subjective emotions for each behavior. However, this perception is immediate and fragmented. Every single occurrence and emotion does not directly shape real humans; instead, information is filtered, holistically evaluated and summarized through unique cognitive processes. Instead, humans will filter and evaluate the information we receive. Such a process is Interesting; psychologists also apply computational thinking to understand how humans absorb, comprehend information, and shape their personalities, viewing the human brain as an information processing system that identifies, stores and ultimately reorganizes and reconstructs its system(solso_cognitive_2005, ). Based on this concept, we developed the Memory and Insight processes within the Cognition module.
@@
Memory - Characteristic Information Filter: Memory serves as a universal module in the current Agent design, with the Generative Agent architecture introducing the Memory-stream architecture for storing, encoding, and retrieving relevant experiences. Cognitive psychology suggests that people with different personalities tend to notice and acquire information about the objective world differently(minsky_society_1988, )(basu_embodied_2004, ), indicating the need for a Memory system tailored to individual traits. Therefore, the Evolving Agents place greater emphasis on differentiated information weight assessment. It includes two types of memory data structures: short-term and long-term memory. Short-term memory stores all executed actions and their corresponding emotions as temporary data. Long-term memory, on the other hand, archives information filtered and summarized from short-term memory. The Agent selects from its existing short-term memory based on the personality features within its Character Structure. It describes multiple behaviors and emotions subjectively, storing the outcomes in its long-term memory database. In practice, Agents develop long-term memories that align with personal traits, especially personality contradictions. For example, Benjamin, contemplating interdisciplinary development, pays attention to information outside his field of study, such as reading a literature book in the library. Isabella, grappling with personal versus societal conflicts, takes great interest in the social philosophy discussions provided by Sophia.
@@
Cognitive psychology’s memory processing theory(solso_cognitive_2005, ) suggests that even long-term memories fade over time, understood as the gradual blurring of information. The design of long-term memory also incorporates this fading mechanism by setting a capacity limit for the memory pool. Upon reaching this threshold, the Agent abstracts and condenses the earliest memories to blur them, as shown in Figure 10.
@@
Figure 10. The long-term memory mechanism of characters involves extracting information of interest from short-term memory and blurring older information.
@@
Insight - Overall Process of the Contextual Incident: Cognitive psychology emphasizes the holism of cognition, where people combine information and relevant contexts to complete an integrated cognitive process(solso_cognitive_2005, ). However, in our system, the information filtered through the memory module remains a vast amount of fragmented data, making it difficult for agents to integrate them for holistic thinking. Therefore, we designed the Insight submodule within the Cognition module to help simulate this holistic thinking process. Moreover, we aim for insight to reflect the diversity of individual cognitive processes. Thus, we enable insight to reference the complete character structure, summarizing and reflecting on all events of the day and the long-term memory. In practical tests, we can intuitively perceive that each Agent has unique insights into daily events.
@@
For example, on the first day, Benjamin realizes that becoming a well-rounded individual requires technical breakthroughs and efforts in humanitarian care and social skills. On the second day, Benjamin sees the potential for integrating technology with the humanities, beginning to incorporate a humanitarian perspective into his technical explorations, seeking a harmonious balance between the two. The procedure of such is shown in Figure 11.
@@ 3.2.4. Character Growth
Chain-like Update Process: The foundational materials for Agent evolution are prepared after processing external events through the Emotion and Cognition modules. The Character Growth module is responsible for simulating the personality update of agents, transforming daily evolutions into comprehensive descriptions updated within the Agent’s Character Structure. This structure doesn’t involve complex algorithmic logic but demands high precision in prompt engineering to depict personality changes accurately. We designed the Character Growth process in a chain-like pattern after extensive testing and consideration of the interrelations within the Character Structure, dividing it into four sequential sub-processes based on the Character Structure dimensions, excluding Basic Information: State Modify, Feature Modify, Contradiction Modify, and Preference Modify. In the design of specific prompts, we maintain Character Structure dimensions, continuously summarizing new personality information based on the Agent’s current state and filling it into the corresponding dimensions. In the continuous assessment of subsequent development, this growth mechanism has provided us with traceable and observable descriptions of personality changes.
@@ 4. Simulation Platform
Like the sandbox simulation environment proposed in the former work, we create a Sandbox using the Unity engine, as shown in Figure 12. In this sandbox, Agents will move and exhibit their actions based on the computational results of Evolving Agents architecture, making it easier for observers to interact.
Figure 12. In the sandbox world, users can observe the Agent’s daily behavioural activities and inner thoughts.
@@ 4.1.1. Three Level Based Environment
Describing information contained in the environment to the large language model is vital. Agents need to be aware of the objects in the environment, whether they are interactable and how to interact. When describing information contained in the environment, we offer a novel thought. We designed a building-place-goal three-level data structure, as shown in Figure 15, to describe the environment. This structure enables us to efficiently convey structured environmental information to large language models, thus allowing agents to fully consider existing environmental information and locations during their planning phase and improving their efficiency in understanding the environment. This directly manifests that large language models will aim to align their planning objectives as closely as possible with the existing environmental functions and descriptions.
@@ 4.1.2. NPC Action Execute
The daily plans generated by the Agent planning module are read and translated into ScriptableObject-type task description files in Unity. When the game time reaches the next plan, the engine will drive the Agent to move to the target location using the A-star algorithm and, upon arrival, play the corresponding animation. To enhance the decoupling of work between the agent architecture and the sandbox, we have detailed programmatic descriptions of the operational rules in the sandbox, such as the Agent’s moving speed and spatial information descriptions. This way, the agent architecture can directly generate Agent performance data that fits the logic of the sandbox (for example, if the Sandbox screen running time from the dormitory to the college is 15 minutes, then the agent architecture will consider this 15-minute moving time in advance when executing). This approach ensures the decoupling level during the team development process and enables the users to design and iterate environment settings when art resources for the sandbox are unavailable.
@@ 4.2.1. Environment Editing
Inspired by behavioral psychology, Evolving Agents is also envisioned as a research tool for the field. In behavioral psychology, external stimuli are one of the primary variables in personality development or change. In this potential demand scenario, we hope the system will allow users to easily modify the system’s environmental settings to meet their future needs. Environmental updates in the game interface require corresponding art resources, but expecting users to provide these resources is impractical. However, as mentioned in section 4.2, agent architecture and sandbox are highly decoupled. Users can edit the environment and obtain simulation results without waiting for the art scenes to update. We have developed a tabular operation experience for environmental editing to lower the learning threshold for user editing. Users can directly control environmental variables by adjusting an online CSV table, thereby achieving adjustments to the Agent’s living space.
Figure 14. Users can directly update environmental settings by adjusting an online CSV table
@@ 4.2.2. Chat With Agent
Although the current research theme does not encourage users to provide artificial stimuli beyond the simulation process for Agents, we also build relevant user interaction experiences in the Simulative System. We aim for user interactions to be consistent with other Agents from an Agent’s perspective. Therefore, users can choose to initiate a conversation with any Agent not currently engaged in a dialogue during the system’s operation. This conversation will also be recorded and summarized in the Agent’s Dialog-memory, contributing to the Agent’s cognitive process.
@@ 5. OBJECTIVE DATA VERIFICATION
Evaluating believability and permeability mentioned above relies on human users’ observation and subjective assessment. However, before recruiting users for experimentation, we can first analyze the operational logs of the Agents to verify changes in personality and behavior patterns that have occurred during the simulation from a machine metrics standpoint (like alpha and beta testing in software development processes).
We implemented two types of objective data validations, using the Big Five personality assessment and behavior difference analysis, respectively. By comparing with the ablation architecture, we objectively verified that the Evolving Agents exhibited personality evolution and corresponding behavioral changes during the simulation process.
@@ 5.1. Big Five Personality Traits Assessment
We aim for Agents to define their current personality based on objective scale questionnaires throughout the simulation process and find the objective evidence of personality changes in Evolving Agents by comparing results over time. The character structure of Evolving Agents is represented based on the Big Five personality traits. Therefore, we refer to existing Big Five personality analysis questionnaires and equip Evolving Agents with the ability to fill out questionnaires, enabling them to continuously undertake Big Five Personality Inventory (BFI-44) assessments(john_big_1999, )and to output the analysis results of the Big Five personality traits during the simulation process. In the specific implementation of enabling Evolving Agents to fill out questionnaires, we discover that even when we constrain the Agent to the same state, the results of each BFI-44 assessment vary significantly. The large language models lack a clear standard for reference and comparison. As a result, we design a process for filling out the same set of questions across multiple days in parallel, asking the model to focus on the differences from day to day for comprehensive questionnaire completion, which effectively resolves the issue of lacking a clear standard and reference for questionnaire completion across sessions.
Contrasted with the complete Evolving Agents architecture, we have an ablated architecture that lacks the insight and Character Growth components of the Personality system. This ablation is because this version of the Evolving Agents architecture is somewhat similar to previous Agent architectures, including the basic processes of Plan-Action-Memory-Reflect. Hence, it has a certain referential value.
We use three Agents, Isabella, Benjamin, and Sophia, for a seven-day simulation within a sandbox(park_generative_2023, ), calculating their daily scores across the five dimensions of the Big Five personality traits, as is shown in Figure 16 and Figure 17, with the scoring method referenced. Finally, we determined the difference in values between each dimension over two days for each person. Let
@@
be the set of dimensions evaluated. For each dimension $d\in D$, given a series of scores $s_{d}$ defined as shown below
@@
The Big Five personality change states of Isabella, Benjamin, and Sophia, who have different initial personality settings over seven days, are shown below
@@
We can see that the Agents in the ablated group show no change in their personality assessment results over the seven days, with an average personality difference of zero every two days for all three Agents. In contrast, Agents in the Full Architecture exhibit active personality change states. For Benjamin, the $\Delta_{\text{overall}}$ is 4.37; for Isabella, it is 6.52; and for Sophia, it is 6.77. This data supports our claim to a great extent that the Evolving Agents can lead to personality evolution in Agents during the simulation process.
From the trends in changes across various dimensions, we also discover that Agents with different personalities exhibit distinct patterns of personality evolution. For example, Benjamin, an introverted and stable computer science student, is slowly adjusting his personality toward his target. In contrast, Sophia, a more creative and active writer, experiences rapid changes in neuroticism and extraversion, which intuitively aligns with our understanding of their individual personalities.
However, while displaying divergent changes, the three individuals also show similar development directions. Overall, all three have shown a decrease in neuroticism, possibly due to the positive bias in the content of large language models. This makes it easier for agents to evolve towards more self-consistent and peaceful personality directions.
@@ 5.2. Behavioral Similarity
In the Agent architecture, we limit the goal selection in the planning phase to a finite number of abstract categories. This allows us to conveniently measure the similarity of daily plans by encoding each day’s plan and analyzing them based on the Euclidean distance, thereby providing a measure of behavioral differences between days. Given a dataset representing daily goal counts, the Euclidean distance between the behavior on two different days, day $i$ and day $j$, is defined by the formula:
@@
where $E(i,j)$ represents the Euclidean distance between day $i$ and day $j$, $x_{ik}$ is the count of goal $k$ on day $i$, and $x_{jk}$ is the count of the same goal on day $j$. $N$ is the total number of distinct goals observed over all days. This calculation assumes that for any goal not observed on a given day, its count is zero.
Given a distance matrix $D$ of size $n\times n$, where $D_{ij}$ represents the Euclidean distance between the behavior on days $i$ and $j$, the overall activity level, denoted as $A$, is calculated as the average of all upper triangular elements of $D$ excluding the diagonal. The formula gives this overall activity level:
@@
We use the same ablation control group as in the Big Five personality trait analysis for the same reason. This is the version of Evolving Agents most similar in architectural description to past studies; hence it holds referential value.
We use three Agents, Isabella, Benjamin, and Sophia, for a seven-day simulation within a Sandbox, calculating the behavioral differences between days. The analysis results of behavioral similarity are as follows:
@@
For Isabella, the activity level in the full architecture is 2.293, while the overall activity level in the ablated architecture is 1.416; for Benjamin, the overall activity level in the full architecture is 2.152, while in the ablated architecture, it is 1.742; for Sophia, the overall activity level in the full architecture is 1.949, while in the ablated architecture it is 1.474. It is evident that compared with the ablated Agent architecture, Agents in the full architecture exhibited more active behavioral changes.
When specifically planning behavior in the Plan module, there are no differences in the input data composition between the full architecture and the ablated architecture of Agent planning. The only difference is that the full architecture agents receive evolved personalities daily, whereas the ablated architecture repeats the initial personality settings. This also indirectly verifies that personality differences are one reason for changes in behavior patterns.